A small sphere 1 carrying charge q1 and having inertia m1 is constrained to moving inside a narrow vertical tube Fixed at the bottom of the tube is a small sphere 2 carrying .

.charge q2



Determine the equilibrium height h

for sphere 1 (ignore friction).

Express your answer in terms of Coulomb's constant k

, acceleration due to gravity g

, and the variables q1, q2, and m1






1
Expert's answer
2021-09-04T07:34:44-0400

There are two forces acting on the sphere 1: Coulomb's force acting up and gravitation acting down. The first one at height "h" is given as follows:


"F_C = k\dfrac{q_1q_2}{h^2}"

The second one at any height (small compared with the size of the Earth) is given as follows:


"F_g = m_1g"

At equilibrium position they are equal to each other. Thus, obtain for height:


"F_C = F_g\\""k\dfrac{q_1q_2}{h^2} = m_1g\\
kq_1q_2 = h^2m_1g\\
h = \sqrt{\dfrac{kq_1q_2}{m_1g}}"

Answer. "h = \sqrt{\dfrac{kq_1q_2}{m_1g}}".